OpenStudy (anonymous):

The length of a rectangle is 11 ft less than three times the width, and the area of the rectangle is 70 ft2 . Find the dimensions of the rectangle.

OpenStudy (campbell_st):

well if the width is w then the length is l = 3w - 11 and you know \[70 = w \times (3w - 11)\] so you need to solve the equation \[3w^2 - 11w - 70 = 0\]